OpenStudy (anonymous):

state the property of real numbers or the property of inequality that justifies each step in the solution of the equation 3x+5=8x

OpenStudy (anonymous):

3x + 5 = 8 x 3x + 5 + (–3x) = 8x + (–3x) 3x + [(–3x) + 5] = 8x + (–3x) [3 x + (–3x)] + 5 = 8x + (–3x) 0 + 5 = 8x + (–3x) 5 = 8x + (–3x) 5 = [8 + (–3)]x 5 = 5 • x 15 • 5 = 15 (5x) 15 • 5 = (15 ∙5)x 1 = 1 • x 1 = x x = 1

OpenStudy (anonymous):

There are 4 properties to choose from. Look them up.

OpenStudy (anonymous):

I know the basic properties. I reall just want to make sure I have the correct properties for each step.

OpenStudy (anonymous):

what have you got so far?

OpenStudy (anonymous):

3x+5+(-3x)=8x+(-3x) Distributive; 3x+[(-3x)+5}=8x+(-3x) Commutative; [3x+(-3x)]+5=8x(-3x) Commutative; 0+5=8x+(-3x) Additive Identity; 5=8x+(-3x) Commutative; 5=[8+(-3)]x Order of operations 5=5 * x Substitution; 1/5 * 5=1/5(5x) Multiplicative Inverse Property